# Learning Task-adaptive Quasi-stiffness Control for a Powered Transfemoral Prosthesis

### v1

@@ IV Discussion
IV-A Evaluation of the Proposed Control Method
This work proposed a learning task-adaptive quasi-stiffness controller for a powered transfemoral prosthesis that generalizes across tasks. The proposed method has the potential to quickly and autonomously configure the impedance parameters for variable tasks. Experimental results and walking performance (see the accompanying video) highlight its capacity to replicate able-bodied biomechanics across variable tasks. As shown in Fig. 7, the observed kinematic and kinetic profiles of the powered prosthetic joint vary with walking speed and are similar to able-bodied references, such as appropriately varying peak angles, peak moments, and trends of profiles. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which indicates that the quai-stiffness control has the potential to improve gait symmetry. Knee kinematics show the most differences with respect to able-bodied data, which may be because the torque-angle relationship of the knee joint is more complex with fewer linear parts, which inevitably introduces errors when using linear approximations.
@@ IV-B Limitations and Future Work
A limitation of this study is that the torque-angle relationship of the swing phase tends to be more nonlinear and complex as the task becomes more difficult (higher walking speed or more challenging terrain conditions), which means that using linear regression to estimate quasi-stiffness may introduce larger errors. In future work, we will investigate the use of nonlinear regression to estimate the swing phase quasi-stiffness. Nevertheless, this work focuses on designing a powered transfemoral prosthesis control framework that generalizes across various tasks. Experimental results and walking performance (see the accompanying video) preliminarily verify that the framework has the potential to achieve task-adaptive control. Future studies will invite more participants to verify the performance of the proposed framework in outdoor experiments.
V Conclusion
This study proposed a learning task-adaptive quasi-stiffness control framework for a powered transfemoral prosthesis working over variable tasks. Learning from the human demonstration method is developed for the torque-angle relationship reconstruction in a new task, which consists of a GPR-based target features prediction model and a KMP-based trajectories reconstruction model. A transfemoral amputee participant demonstrated the control framework’s ability to generalize across tasks by varying speeds and inclines walking experiments. The experimental results validate that the proposed control framework has the potential to reproduce biomimetic joint torques and angles for a powered prosthesis during variable-task walking.

### v2

@@ IV Discussion
IV-A Evaluation of the Proposed Framework
This work proposed a learning task-adaptive quasi-stiffness controller for a powered transfemoral prosthesis that generalizes across tasks. The proposed method has the potential to quickly and autonomously configure the impedance parameters for variable tasks. Experimental results and walking performance (see the accompanying video) highlight its capacity to replicate able-bodied biomechanics across variable tasks. As shown in Fig. 7, the observed kinematic and kinetic profiles of the powered prosthetic joint vary with walking tasks and are similar to able-bodied references, such as appropriately varying peak angles, peak moments, and trends of profiles. As speed or incline increases, the proposed method can generate biomechanical performances similar to the intact limb of the user and healthy individuals. For instance, as the incline increases, ankle joint torque correspondingly increases to provide individuals with limb amputations more push-off torque. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories (also the lower RMSE of the observed kinematics with respect to the intact leg kinematics shown in Fig. 8(a)), which indicates that the quasi-stiffness control has the potential to improve gait symmetry. Knee kinematics show the most differences with respect to able-bodied data, which may be because the torque-angle relationship of the knee joint is more complex with fewer linear parts, which inevitably introduces errors when using linear approximations.
This article introduces the widely used FSMIC controller as a benchmark for comparison. Fig. 9 and 10 show the proposed learning quasi-stiffness control framework meets the benchmark’s performance in most walking tasks. Fig. 10(b) indicates that the proposed method surpasses the benchmark’s performance in generating the biological kinetics. As shown in Fig. 9 and Fig. 10(a), the FSMIC’s performance was not drastically worse than the proposed method in the tested metrics, it required an average of 20 min of tuning per tuned task. The proposed method highlights its advantages in reproducing biomimetic kinematics and kinetics similar to able-bodied references, aligns with the performance of the FSMIC without requiring manual impedance tuning, and reveals the potential to enable autonomous task adaptation during continuously varying walking tasks.
@@ IV-B Limitations and Future Work
A limitation of this study is that the torque-angle relationship of the swing phase tends to be more nonlinear and complex as the task becomes more difficult (composite tasks, such as walking speed and incline changing simultaneously), which means that using linear regression to estimate quasi-stiffness may introduce more significant errors. In future work, we will investigate using nonlinear regression to estimate the swing phase quasi-stiffness. Nevertheless, this work focuses on designing a powered transfemoral prosthesis control framework that generalizes across various tasks. Experimental results and walking performance (see the accompanying video) preliminarily verify that the framework has the potential to achieve task-adaptive control. Future studies will focus on evaluating the performance of the proposed method in stair climbing tasks. Besides, we will invite more participants to verify the performance of the proposed framework and investigate the user preference of the proposed method and the FSMIC in different tasks.